A ball tossed in the air vertically from ground level returns to earth 4 seconds later. Find the inital velocity and maximum height of the ball.


1
Expert's answer
2021-10-14T07:47:26-0400

According to the Galilei formulas:

"s(t) =s{\scriptscriptstyle 0}+v{\scriptscriptstyle 0}*t - {\frac 1 2}gt^{2}"

"v(t) = v{\scriptscriptstyle 0}-gt"

Where "s{\scriptscriptstyle 0}" - initial height, "v{\scriptscriptstyle 0}" - initial speed, g - acceleration of the gravity(9.8 m/s), t - point of time

"s{\scriptscriptstyle 0} = 0 \to s(t) = v{\scriptscriptstyle 0}*t - {\frac 1 2}gt^{2}"

The graph of s(t) takes form of parabola with branches pointing down(since "-{\frac 1 2}g <0") , which means the t-point of the vertex of the parabola is the point where value of s(t) is maximized(maximum height of the ball)

We have s(0) = s(4), which means points t = 0 and t = 4 are located on the different branch and equidistant from the axis of the symmetry of parabola(which takes form x = "t{\scriptscriptstyle m}" , where "t{\scriptscriptstyle m}" - t-point of the vertex of the parabola.

So, "t{\scriptscriptstyle m} = {\frac {0+4} 2} = 2"

It took 2 seconds for the ball to reach it's maximum height

"s(2) =2v{\scriptscriptstyle 0} - 2g"

To find the value of s(2) we should find "v{\scriptscriptstyle 0}"

Since s(2) is maximum, then "s'(2) = v(2) =0"

"v(2) = v{\scriptscriptstyle 0}-2g"

"v{\scriptscriptstyle 0} = 2g"

"v{\scriptscriptstyle 0} = 19.6" m/s

So, by substituting the obtained values we get

"s(2) =2*19.6 - 2*9.8 = 19.6" m

The initial velocity was 19.6 m/s, the maximum height was 19.6 m.